OpenStudy (anonymous):

A bathtub contains 20 gallons of water. After the drain is opened, there are 14 gallons of water remaining after 30 seconds. Using this information, how can you describe the rate of change in the amount of water remaining in the tub per second? a decrease of 0.5 gallons per second an increase of 0.5 gallons per second a decrease of 0.2 gallons per second an increase of 0.2 gallons per second

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@UnkleRhaukus @imqwerty

imqwerty (imqwerty):

rate of change in amount of water= change in volume/time taken rate of change in amount of water=(20-14)/30 rate of change in amount of water=6/30 rate of change in amount of water=1/5 rate of change in amount of water=0.2gallons per second

imqwerty (imqwerty):

yes because the volume is decreasing so we will say a decrease of 0.2 gallons per second

OpenStudy (anonymous):

a decrease of 0.2 gallons per second?

OpenStudy (anonymous):

ok thanks!

hartnn (hartnn):

or mathematically, (final amount - initial amount)/ time = (14-20)/30 = - 0.2 this negative suggest that it will be decrease of 0.2 gallons per sec :)

imqwerty (imqwerty):

np